Head-First Company now sells both bicycle helmets and motorcycle helmets. Next year, Head- First expects to produce total revenue of $575,000 and incur total variable cost of $400,000. Total fixed cost is expected to be $60,500.

Required:
1. Calculate the break-even point in sales dollars for Head-First. Round the contribution margin ratio to four decimal places and sales to the nearest dollar.
2. Check your answer by preparing a contribution margin income statement.

1. Calculate the break-even point in sales dollars for Head-First. Round the contribution margin ratio to four decimal places and sales to the nearest dollar.

The break-even point in sales equals $
Answer #1

Sales = $575,000

Variable cost = $400,000

Contribution margin = Sales - Variable cost

= 575,000-400,000

= $175,000

Contribution margin ratio = Contribution margin/Sales

= 175,000/575,000

= 30.4348%

Break-even point in sales dollars = Fixed cost / Contribution margin ratio

= 60,500/30.4348%

= $198,786

2.

Contribution Margin Income Statement
Sales 198,786
Variable cost { 198,786 x (100% - 30.4348%)} -138,286
Contribution margin 60,500
Fixed cost -60,500
Net operating income $0

Kindly comment if you need further assistance. Thanks